OpenStudy (anonymous):

Factor completely: 2x2 − 2x − 40 2(x − 5)(x + 4) (2x − 10)(x + 4) (x − 5)(2x + 8) 2(x − 4)(x + 5)

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (anonymous):

B?

jigglypuff314 (jigglypuff314):

factor out 2 so 2x^2 − 2x − 40 = 2(x^2 - x - 20) ^ ^ so should multiply to -20 and add to -1 and should have 2 in front because factored

OpenStudy (anonymous):

so C?

jigglypuff314 (jigglypuff314):

does C have a 2 in front?

OpenStudy (anonymous):

Oops I meant D.

jigglypuff314 (jigglypuff314):

2(x − 4)(x + 5) ^ ^ does -4 + 5 = -1 ?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

so it's A?

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

Wooo whooo! haha